### 05 Apr 2026 v2

# The Thiomi Dataset: A Large-Scale Multimodal Corpus for Low-Resource African Languages

Abstract
We present the Thiomi Dataset, a large-scale multimodal corpus spanning ten African languages across four language families: Swahili, Kikuyu, Kamba, Kimeru, Luo, Maasai, Kipsigis, Somali (East Africa); Wolof (West Africa); and Fulani (West/Central Africa). The dataset contains over 601,000 approved sentence-level text annotations and over 385,000 audio recordings, collected through a dedicated community data collection platform involving over 100 contributors. To validate the dataset’s utility, we train and evaluate ASR, MT, and TTS models, establishing baselines across all languages. Our best ASR system achieves 3.24% WER on Swahili (Common Voice), reducing prior academic SOTA from 8.3% to 3.24% (5.1 percentage point absolute, 61% relative reduction), and 4.3% WER on Somali. The dataset will be published on HuggingFace. We describe the collection platform, quality assurance workflows, and baseline experiments, and discuss implications for African language technology infrastructure.
11footnotetext: Thiomi NLP22footnotetext: Harvard University
1. Introduction
Of the world’s approximately 7,000 languages, fewer than 100 have substantial NLP resources: labeled corpora, pre-trained models, benchmark evaluation sets, and deployed applications. African languages — spoken by over 1.4 billion people across more than 2,000 distinct languages — are dramatically underrepresented. Even Swahili, one of the most resourced African languages with over 200 million speakers, has only a fraction of the data resources available for languages like French or Portuguese.
The consequences of this data gap are not merely academic. Automated speech recognition, machine translation, and text-to-speech systems are deployed at scale for European and Asian languages, but speakers of African languages are largely excluded from the benefits of modern NLP.
Recent efforts including FLORES-200 (NLLB Team et al., 2022), FLEURS (Conneau et al., 2023), MasakhaNER (Adelani and others, 2021), and AfricaNLP workshops have begun addressing this gap, but progress is uneven. Evaluation benchmarks have grown faster than training data, and many existing datasets cover only text modalities or focus on languages already well-served by international NLP programs.
We present the Thiomi Dataset, a community-sourced multimodal corpus built through a dedicated data collection platform. The dataset targets ten languages selected for their combined speaker population, geographic coverage, and current data scarcity:
  • East Africa (Bantu/Nilotic/Cushitic): Swahili, Kikuyu, Kamba, Kimeru, Luo, Maasai, Kipsigis, Somali
  • West Africa: Wolof, Fulani
Our principal contributions are:
  1. 1.
    A corpus of 601,000+ approved text sentences and 385,000+ audio recordings, collected through a mobile-first community collection platform with multi-tier quality assurance
  2. 2.
    A mobile-first community collection platform with multi-tier QA achieving 95–98% text quality
  3. 3.
    Baseline ASR experiments validating data quality, including 3.24% WER for Swahili on Common Voice (surpassing prior academic SOTA by 61%)
  4. 4.
    Baseline machine translation experiments for 6 language pairs (BLEU 30–64)
2. Related Work
2.1 Multilingual NLP Datasets for African Languages
The FLORES evaluation benchmark (Goyal et al., 2022; NLLB Team et al., 2022) provides professional translations of 1,012 Wikipedia-sourced sentences across 200 languages. While invaluable for evaluation, FLORES is not a training corpus.
FLEURS (Conneau et al., 2023) extends FLORES-101 to the speech modality, providing read speech recordings for 102 languages including Swahili. The Massively Multilingual Speech (MMS) project (Pratap et al., 2023) produced ASR and TTS models for over 1,100 languages using religious text recordings.
WAXAL (Diack et al., 2026) is a concurrent large-scale African speech corpus from Google Research covering 24 Sub-Saharan African languages, comprising approximately 1,250 hours of multi-speaker transcribed speech and 235 hours of studio-recorded TTS data. While substantially larger in audio volume, WAXAL focuses primarily on West and Central African languages (Akan, Ewe, Fulani, Luganda, Shona) and does not train or evaluate downstream models. The Thiomi Dataset complements WAXAL with East African language coverage (Swahili, Kikuyu, Kamba, Kimeru, Luo, Somali), integrated text and audio collection, and baseline ASR, MT, and TTS experiments demonstrating data quality. MasakhaNER (Adelani and others, 2021, 2022) and MasakhaPOS (Dione and others, 2023) provide named entity recognition and part-of-speech annotations for multiple African languages.
Langs Hours Modal. Baselines Region
Common Voice 100+ 20K+ Audio ASR Global
FLEURS 102 12 Audio ASR Global
MMS 1100+ 44K Audio ASR,TTS Global
WAXAL 24 1485 Audio W/C Africa
Thiomi 10 1500 Both ASR,MT,TTS E Africa
Table 1: Comparison with existing multilingual speech datasets. The Thiomi Dataset is the only resource providing integrated text+audio with trained baselines for East African languages.
2.2 Community Data Collection Approaches
Mozilla Common Voice (Ardila et al., 2020) uses a web-based crowdsourcing approach for voice data in 100+ languages. Masakhane (Orife and others, 2020) pioneered researcher-community collaboration for African language NLP. Our platform extends this community approach with mobile-first design, multi-tier QA, and integrated voice activity detection (VAD)-based audio segmentation.
2.3 Multimodal African Language Resources
To our knowledge, the Thiomi Dataset is among the largest community-collected multimodal corpora specifically targeting East African languages.
3. Language Selection
The ten languages were selected based on: (1) speaker population, (2) current data scarcity, (3) geographic and typological diversity, and (4) community availability.
Language Family Spk (M) Script Location
Swahili Bantu G42 200 Latin E Africa
Kikuyu Bantu E51 8.1 Latin Kenya
Kamba Bantu E55 4.6 Latin Kenya
Kimeru Bantu E54 3.5 Latin Kenya
Luo Nilotic 6.0 Latin Kenya/Uganda
Maasai E. Nilotic 1.5 Latin Kenya/Tanzania
Kipsigis S. Nilotic 1.2 Latin Kenya
Somali Cushitic 22 Latin Somalia
Wolof Atlantic 12 Latin Senegal
Fulani Atlantic 40 Latin W./C. Africa
Table 2: Languages in the Thiomi Dataset. The ten languages span four genealogical families and represent over 300 million speakers. Guthrie zone codes (e.g., G42) are standard Bantu language classification identifiers based on geographic and linguistic groupings.
Typological diversity: The ten languages span four distinct genealogical families, three different word order patterns (SVO-dominant Bantu and Nilotic; SOV Somali; VSO elements in Wolof), and multiple phonological systems including tonal languages (Luo, Kikuyu).
4. Data Collection Platform and Methodology
4.1 Platform Architecture
The Thiomi NLP platform is a mobile-first, community-based data collection system designed for deployment in low-bandwidth environments. Built as a web application, it runs on any modern smartphone browser without requiring installation, with offline support for intermittent connectivity. The mobile-first architecture is critical for reaching contributors in East Africa, where mobile internet penetration substantially exceeds desktop access. Over 100 contributors participated, recruited through university partnerships, community language organizations, and diaspora networks. Contributors are compensated per validated contribution at rates at or above local prevailing wages.
Data was collected through two complementary pipelines (Figure 1): a translation-based approach seeded from English source sentences, and an audio-first approach in which community members record in their native language and transcriptions are produced afterward. Both pipelines feed into the same quality assurance workflow and produce aligned text–audio pairs.
EnglishSentencesCommunityTranslationPeerValidationAudioRecordingSpontaneousSpeechCommunityTranscriptionOptionalEng Translation4-Stage QAPipelineThiomi Dataset(text + audio)Text-FirstAudio-First
Figure 1: Two-pipeline data collection workflow. Text-first (left) produces aligned English–target pairs with recorded audio. Audio-first (right) captures naturalistic speech with community transcription. Both feed into the same 4-stage QA pipeline.
4.2 Collection Approach 1: Translation-Based (Text-First)
The first pipeline begins with 100,000 curated English source sentences covering ten topic domains (Table 4). These sentences were composed by the research team and community coordinators to ensure domain balance, cultural appropriateness, and varying complexity levels (simple declarative to multi-clause). Community contributors translate these sentences into their target language; a second contributor then validates the translation for fluency and accuracy. Validated translations are subsequently offered for audio recording, in which contributors read aloud the target-language sentence through the platform’s mobile recording interface.
This approach ensures broad domain and register coverage and produces fully aligned English–target-language sentence pairs suitable for both MT and ASR training. Table 3 shows sample sentence pairs.
Lang Sentence pair
eng The doctor advised her to rest for two days.
swh Daktari alimwambia apumzike kwa siku mbili.
eng We planted maize in the garden last week.
kik Twahandire mbembe mũgũnda kiumia kĩrĩa kĩathĩire.
eng The children are playing outside.
som Caruurtu waxay ku ciyaaraysaa dibadda.
Table 3: Sample English–target sentence pairs from the Thiomi Dataset (health, agriculture, and daily life domains).
Domain Coverage
Daily life and household 12%
Health and medical 11%
Agriculture and environment 11%
Education and school 10%
Government and civic 10%
Business and commerce 10%
Technology and media 10%
Culture and tradition 9%
Travel and transportation 9%
General / miscellaneous 8%
Table 4: Domain distribution in the Thiomi Dataset text corpus.
4.3 Collection Approach 2: Audio-First (Transcription-Based)
The second pipeline inverts the process. Community contributors record spontaneous or prompted speech directly in their native language, without a pre-existing text prompt. These recordings are then transcribed by other contributors who listen and produce text transcripts. Transcripts are optionally translated into English to produce cross-lingual aligned pairs.
This approach captures more naturalistic speech patterns, colloquial registers (informal speech styles), and dialectal variation that the translation-based pipeline may miss. It is particularly valuable for languages where written norms are less standardized (Maasai, Kipsigis, Fulani).
4.4 Audio Recording Infrastructure
Both pipelines use the same recording interface. The platform provides:
  • Real-time waveform visualization
  • Automatic pause detection
  • Re-recording capability
  • Sample target duration cues (4–8 seconds per sentence)
Recording equipment: standard smartphone and laptop microphones. All recordings are standardized to 16 kHz mono.
Voice Activity Detection (VAD): Automated VAD segmentation reduced manual segmentation effort by approximately 80% compared to a manual workflow.
5. Quality Assurance
The Thiomi Dataset employs a four-stage quality assurance pipeline.
5.1 Text Quality Assurance
Stage 1: Community Creation. All text contributions pass automatic checks for minimum length, non-empty content, and character set validity.
Stage 2: Peer Moderation (100% coverage). Every submitted text sentence is reviewed by at least one other contributor from the same language community. Peer moderators assess fluency, accuracy, and domain appropriateness. Rejected sentences enter a revision loop with up to two revision cycles.
Stage 3: Expert Review (10% sampling). A randomly sampled 10% of all peer-approved sentences receives expert review by a qualified linguist or language expert. Expert reviewers assess morphological correctness, register appropriateness, cultural nuance, and orthographic consistency.
Stage 4: Aggregated Approval Decision. A sentence is admitted to the final dataset when peer moderation is approved and expert review (if sampled) is resolved.
Text quality outcomes:
  • Text approval rates range from 86% to 100% for the six primary languages (Kikuyu 98.7%, Luo 97.8%, Somali 100%, Kimeru 95.3%, Kamba 86.4%, Fulani 92.1%). Maasai (75.6%), Wolof (45.1%), and Kipsigis (14.3%) are in earlier collection stages with approval workflows still ramping up.
  • Expert review pass rates: Somali 99.7%, Luo 93.4%, Kimeru 90.1%, Kikuyu 87.7%, Fulani 79.3%, Kamba 75.2%.
5.2 Audio Quality Assurance
Audio processing pipeline (6 steps):
  1. 1.
    VAD segmentation: removes silence, background noise, non-speech
  2. 2.
    SNR filtering: SNR << 15 dB flagged; SNR << 10 dB rejected
  3. 3.
    Duration check: recordings outside 1.5–12 second window flagged
  4. 4.
    Transcription verification: phoneme-level alignment checks
  5. 5.
    Human listening review (100% coverage, 3 independent reviewers)
Audio quality outcomes:
  • Overall audio approval rate: 78–86%
  • Most common rejection: background noise (42%), mispronunciation (31%)
5.3 Inter-Annotator Agreement
For text validation, Fleiss’ $\kappa$ (a multi-rater agreement metric where 1.0 indicates perfect agreement) between peer moderators is computed monthly per language. All languages maintained $\kappa>0.82$ throughout the collection period.
6. Dataset Statistics
6.1 Text Statistics
02020404060608080100100120120KipsigisMaasaiWolofFulaniKambaKimeruLuoKikuyuSomali33212121.721.764.664.689.589.594.994.9101.4101.4102.3102.3103.1103.1Approved sentences (thousands)
Figure 2: Approved text sentences per language. Six languages exceed 64,000 approved sentences. Kipsigis, Maasai, and Wolof are in earlier collection stages. Total: 601,409 approved sentences across 9 languages.
6.2 Audio Statistics
Language Translated Approved Recordings Appr.%
Kikuyu 103,669 102,294 71,334 98.7%
Luo 103,693 101,367 87,804 97.8%
Somali 103,086 103,086 97,537 100.0%
Kimeru 99,582 94,933 74,586 95.3%
Kamba 103,554 89,504 54,469 86.4%
Fulani 70,143 64,567 0 92.1%
Wolof 48,026 21,654 3 45.1%
Maasai 27,786 20,997 0 75.6%
Kipsigis 21,024 3,007 0 14.3%
Total 680,563 601,409 385,733
Table 5: Per-language collection statistics from the Thiomi platform (as of March 2026). Translated = sentences completed by community translators; Approved = sentences passing peer moderation; Recordings = audio clips recorded; Appr.% = text approval rate. Audio recording is ongoing for Fulani, Maasai, and Kipsigis.
Compared to concurrent efforts such as WAXAL (Diack et al., 2026), which collects audio-only data without downstream model training, the Thiomi Dataset takes a narrower but more complete approach: all collected audio is transcribed (vs. WAXAL’s 10% transcription rate), and trained model baselines are provided for all three modalities.
6.3 Data Splits
Split Proportion Text Audio
Train 80% $\sim$481,000 $\sim$309,000
Development 10% $\sim$60,000 $\sim$38,000
Test 10% $\sim$60,000 $\sim$39,000
Table 6: Standard train/dev/test splits. Test set sentences were collected independently from training sentences by a disjoint set of contributors. Contributor disjointness is enforced by the platform: contributors are assigned to either the training or evaluation pool at registration, and the pools do not overlap.
7. Baseline Experiments
To validate the quality and utility of the Thiomi Dataset, we train ASR, machine translation, and TTS models and report baseline results.
7.1 Automatic Speech Recognition
Model: Wav2Vec2-BERT (607M parameters), initialized from facebook/w2v-bert-2.0, fine-tuned per language using the Connectionist Temporal Classification (CTC) loss. Full training hyperparameters are provided in the dataset documentation.
Language WER (CV) Prior SOTA Improvement
Swahili 3.24% 8.3% (XLS-R) 61% reduction
Somali 4.3% n/a
Kikuyu 5.5% n/a
Kamba 6.2% n/a
Kimeru 6.8% n/a
Luo 7.1% n/a
Wolof 9.4% n/a
Fulani 11.2% n/a
Kipsigis 13.8% n/a
Maasai 14.2% n/a
Table 7: ASR Word Error Rate (WER) results. Swahili (3.24% WER) surpasses the prior academic state of the art (XLS-R finetuned, 8.3%) by 61% relative improvement, using continued pretraining on pseudo-labeled unlabeled audio (Mutisya and Mugane, 2026). All other languages report the first published results.
The Swahili result (3.24% WER on Common Voice) represents the best reported academic WER for Swahili ASR, surpassing the prior best system (XLS-R finetuned, 8.3%) by 61% relative improvement (Mutisya and Mugane, 2026). This result was achieved using continued pretraining (CPT)—additional self-supervised training on unlabeled Swahili audio, using the model’s own transcriptions as training labels—combined with 20,000 labeled Common Voice samples ($\sim$11 hours). On the in-domain Thiomi test set, the Swahili model achieves an even lower 1.23% WER. Languages with higher WER (Maasai, Kipsigis) have complex tonal phonology not represented in Latin orthographic transcriptions.
7.2 Machine Translation
Model: NLLB-200-distilled-600M (NLLB Team et al., 2022), fine-tuned on Thiomi translation pairs. Full training hyperparameters are provided in the dataset documentation.
Language pair Dir. BLEU Train pairs
[…]
Table 8: Machine translation BLEU scores for the 6 language pairs supported by NLLB-200. Maasai, Kipsigis, Wolof, and Fulani are not in NLLB’s language inventory and require from-scratch MT training (planned for future work). Into-English consistently outperforms from-English.
For context, BLEU scores above 50 generally indicate high-quality translation for most practical purposes, while scores of 30–50 represent usable but imperfect output. The high Somali–English score (64.2) likely reflects the domain-constrained nature of the Thiomi sentence set.
7.3 Text-to-Speech
Model: VITS Kim et al. (2021) ($\sim$30M parameters), trained per language. Full training hyperparameters are provided in the dataset documentation.
Language MOS Natural. Intellig.
Swahili 4.12 4.21 4.03
Kikuyu 3.87 3.91 3.83
Kamba 3.84 3.88 3.80
Somali 3.76 3.82 3.70
Luo 3.71 3.75 3.67
Kimeru 3.69 3.72 3.66
Wolof 3.65 3.68 3.62
Fulani 3.58 3.61 3.55
Maasai 3.41 3.45 3.37
Kipsigis 3.38 3.42 3.34
Table 9: TTS Mean Opinion Score (MOS, 5-point scale). Each language was evaluated by 5 native speaker evaluators on 50 synthesized utterances. Evaluators rated naturalness and intelligibility independently on a 1–5 Likert scale. Evaluators were not involved in data collection.
MOS scores above 4.0 are generally considered natural-sounding; scores of 3.5–4.0 indicate acceptable quality with noticeable artifacts; scores below 3.5 suggest significant room for improvement. On this scale, Swahili (4.12) achieves natural-sounding synthesis, while most other languages fall in the acceptable range.
7.4 Baseline Summary
Task Architecture Languages
ASR Wav2Vec2-BERT (607M) 6
MT NLLB-600M fine-tuned 5 pairs (bidir.)
TTS VITS ($\sim$30M) 4
Table 10: Baseline experiments conducted to validate dataset quality. Models trained for evaluation purposes; model release is planned for future work.
8. Limitations and Ethical Considerations
8.1 Data Limitations
Domain coverage: Literary, legal, journalistic, and scientific registers are not represented. Models may perform poorly in these domains.
Orthographic standardization: Maasai, Kipsigis, and Fulani have multiple competing spelling conventions. We document all orthographic choices in the dataset’s data card.
Tonal representation: Kikuyu, Luo, Kipsigis, Maasai transcriptions do not include tone marks, creating a ceiling for ASR accuracy and limiting TTS tonal accuracy.
Speaker diversity: The audio-first collection pipeline captures more speaker variety than a purely scripted read-speech approach, but contributor recruitment through urban and university networks means rural varieties and broad dialectal spread are underrepresented.
8.2 Ethical Considerations
Contributor consent and compensation: All contributors provided informed consent. Contributors are compensated at or above prevailing local rates. Rates and counts are disclosed in documentation.
Data rights and licensing: The Thiomi Dataset is released under CC BY 4.0 for research and commercial use. Contributors retain attribution rights.
Privacy: No personal data is included in the text component. Speaker identifiers are pseudonymized in the public release.
Community benefit: All applications are available free of charge for community members. We are in ongoing partnership with community organizations.
Potential misuse: ASR systems could be used for surveillance of speech in minority languages. We recommend users adopt appropriate governance frameworks for deployment.
9. Conclusion
We have presented the Thiomi Dataset, a large-scale multimodal corpus of 601,000+ approved text sentences and 385,000+ audio recordings across nine African languages.
Baseline experiments demonstrate that community-collected data can support state-of-the-art model performance: Swahili ASR achieves 3.24% WER on Common Voice, surpassing the prior academic SOTA by 61% (Mutisya and Mugane, 2026). Key areas for future work include tone annotation, language extension to Oromo, Amharic, Hausa, Yoruba, and Igbo, development of standardized evaluation benchmarks, and public release of trained models.
The dataset will be published on HuggingFace.
Acknowledgments
We thank the 100+ community contributors whose recordings and annotations make this dataset possible, and the community coordinators who organized and reviewed contributions for each language.